OpenStudy (anonymous):

a vertical parabola has vertex at (-3,-2) and passes through the point (-1,7). Find the equation of the parabola, the equation of the directrix, and the coordinates of the focus point. Please explain!!

OpenStudy (anonymous):

let's get the vertex form first and then work the others from there... so the vertex is at (-3, -2) and the parabola passes through the point (-1, 7). we can use this to get the vertex form of the parabola first... \[y=f(x)=a(x-h)^2+k\]this is the vertex form with vertex at (h, k). sub in those know values. what do you get?

OpenStudy (anonymous):

y=a(x+3)^2-2

OpenStudy (anonymous):

good. now use the given point to find what a is.

OpenStudy (anonymous):

9/4

OpenStudy (anonymous):

is that correct

OpenStudy (anonymous):

yeppers! so know to translate to the directrix thingymabobber... \[4p(y-k)=(x-h)^2\] has distance p to the focus or directrix from the vertex (h, k)

OpenStudy (anonymous):

you just found\[y=\frac{5}{4}(x+3)^2-2\]now work it into the form above and find p.

OpenStudy (anonymous):

you mean 9/4?

OpenStudy (anonymous):

yeah, sorry

OpenStudy (anonymous):

p= 1/9? which is the directrix?

OpenStudy (anonymous):

p = 1/9 but i't not the directrix... it's the distance from the vertex to the focus and it's the distance from the vertex to the directrix.

OpenStudy (anonymous):

so how does that help me find the focus and the directrix

OpenStudy (anonymous):

directrix is gonna be like this... y = k - p focus is at the point (h, k + p)

OpenStudy (anonymous):

Thannkss sooo muchhh!!!! do you know multidimensional calculus?

OpenStudy (anonymous):

it's been a while... whatcha got?

OpenStudy (anonymous):

im taking the class and we just calculated the volume of an ellipse using line integrals derived from double integrals. if I wanted to calculate the volume of an ellipsoid, how would do that?

OpenStudy (anonymous):

you'd be better off asking someone more familiar than me. i could do it but it would be long and arduous.

OpenStudy (anonymous):

okay thanks!! I have one more question.. will u be on later